### v3

# DiFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving

Abstract
Current end-to-end autonomous driving methods resort to unifying modular designs for various tasks (e.g. perception, prediction and planning). Although optimized in a planning-oriented spirit with a fully differentiable framework, existing end-to-end driving systems without ego-centric designs still suffer from unsatisfactory performance and inferior efficiency, owing to the rasterized scene representation learning and redundant information transmission. In this paper, we revisit the human driving behavior and propose an ego-centric fully sparse paradigm, named DiFSD, for end-to-end self-driving. Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner. The sparse perception module performs detection, tracking and online mapping based on sparse representation of the driving scene. The hierarchical interaction module aims to select the Closest In-Path Vehicle / Stationary (CIPV / CIPS) from coarse to fine, benefiting from an additional geometric prior. As for the iterative motion planner, both selected interactive agents and ego-vehicle are considered for joint motion prediction, where the output multi-modal ego-trajectories are optimized in an iterative fashion. Besides, both position-level motion diffusion and trajectory-level planning denoising are introduced for uncertainty modeling, thus facilitating the training stability and convergence of the whole framework. Extensive experiments conducted on nuScenes and Bench2Drive datasets demonstrate the superior planning performance and great efficiency of DiFSD.
1 Introduction
Autonomous driving has experienced notable progress in recent years. Traditional driving systems are commonly decoupled into several standalone tasks, e.g. perception, prediction and planning. However, heavily relying on hand-crafted post-processing, the well-established modular systems suffer from information loss and error accumulation across sequential modules. Recently, end-to-end paradigm integrates all tasks into a unified model for planning-oriented optimization, showcasing great potential in pushing the limit of autonomous driving performance.
Literally, existing end-to-end models [11, 41, 17, 36] designed for reliable trajectory planning can be classified into two mainstreams as summarized in Fig. 1(a) and (b). The dense BEV-Centric paradigm [11, 41] performs perception, prediction and planning consecutively upon the shared BEV (Bird’s Eye View) features, which are computationally expensive leading to inferior efficiency. The sparse Query-Centric paradigm [36] utilizes sparse representation to achieve scene understanding and joint motion planning, thus improving the overall efficiency. However, object-intensive motion prediction inevitably causes computational redundancy and violates the driving habits of human drivers, who usually only concentrate on the Closest In-Path Vehicle / Stationary (CIPV / CIPS) which are more likely to affect the driving intention and trajectory planning of ego-vehicle. Meanwhile, excessive interaction with irrelevant agents will be conversely adverse to the ego-planning. Therefore, the planning performance remains unsatisfactory in both planning safety, comfort and personification.
To this end, we propose DiFSD, an Ego-Centric fully sparse paradigm as shown in Fig. 1(c). Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner. In the sparse perception module, multi-scale image features extracted from visual encoder are adopted for object detection, tracking and online mapping simultaneously in a sparse manner. Then the hierarchical interaction performs ego-centric and object-centric dual interaction to select the CIPV / CIPS with the help of an additional geometric prior. Thus the interactive queries can be selected gradually from coarse to fine. As for the motion planner, the mutual information between sparse interactive queries and ego-query is considered for motion prediction in a joint decoder, which is neglected in previous methods [11, 17] but is essential especially in the scenarios like intersections. To ensure the planning rationality and selection accuracy of interactive queries, the iterative planning optimization is further applied to the multi-modal proposal ego-trajectories, through continually updating the reference line and ego-query. Moreover, both position-level motion diffusion and trajectory-level planning denoising are introduced for stable training and fast convergence. It can not only model the uncertain positions of interactive agents for motion prediction, but also enhance the quality of trajectory refinement with arbitrary offsets. With above elaborate designs, DiFSD exhibits the great potential of fully sparse paradigm for end-to-end autonomous driving, which significantly reduces the average L2 error by 56% and collision rate by 92% than UniAD [11] respectively. Notably, our DiFSD-S achieves faster running efficiency as well. In sum, the main contributions of our work are as follows:
We propose an ego-centric Fully Sparse paradigm for end-to-end self-Driving, named as DiFSD, without any computationally intensive dense scene representation learning and redundant environmental modeling, which is proven to be effective and efficient for ego-planning.
We introduce a geometric prior through intention-guided attention, where the Closest In-Path Vehicle / Stationary (CIPV / CIPS) are gradually picked out through ego-centric cross attention and selection. Besides, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are adopted for uncertainty modeling of motion planning respectively.
2 Related Work
2.1 End-to-End Perception
Recent years witness remarkable progress achieved in multi-view 3D detection, which mainly build elaborate designs upon the dense BEV (Bird’s Eye View) or sparse query features. To generate BEV features, LSS [33] lifts 2D image features to 3D space using depth estimation results, which are then splatted into BEV plane. Follow-up works apply such operation to perform view transform for 3D detection task [14, 13, 19, 7]. Differently, some works [20, 40, 12] project a series of predefined BEV queries in 3D space to the image domain for feature sampling. As for the sparse fashion, current methods [38, 26, 27, 25, 24] adopt a set of sparse queries to integrate spatial-temporal aggregations from multi-view image feature sequence for iterative anchor refinement.
Besides, Multi-Object Tracking (MOT) across multi-cameras is also required for downstream tasks. Recent works [44, 46, 42, 32, 35] seek to explore the joint detection and tracking methods by introducing track queries to detect the unique instances continuously. Sparse4Dv3 [24] proposes an advanced 3D detector which takes the spatial-temporal information to propagate temporal instances for identity reserving, thus achieving superior end-to-end performance without additional tracking designs.
2.2 Online Mapping
Maps could provide important static scenario information to ensure driving safety. Current works [18, 28, 23, 43] manage to construct online maps with on-board sensors, instead of using HD-Map which is labor intensive and expensive. HDMapNet [18] achieves this aim through BEV semantic segmentation and heuristic post-processing to generate map instances. VectorMapNet [28] introduces a two-stage auto-regressive transformer to refine map elements consecutively. MapTR [23] regards map elements as a set of points with equivalent permutations, while StreamMapNet [43] adopts a temporal fusion strategy to enhance the performance. However, all of them reply on dense BEV features for online map construction, which is computationally intensive and not extensible to the sparse manner.
2.3 End-to-End Motion Prediction
Motion prediction of surrounding agents in an end-to-end fashion can relieve the accumulative error between standalone models. FaF [31] predicts both current and future bounding boxes from images using a single convolution network. IntentNet [2] attempts to reason high-level behavior and long-term trajectories simultaneously. PnPNet [22] aggregate trajectory-level features for motion prediction through an online tracking module. ViP3D [6] takes images and HD-Map as input, and adopts agent queries to conduct tracking and prediction. PIP [16] further proposes to replace HD-Map with local vectorized map.
2.4 End-to-End Planning
End-to-end planning paradigm either unites modules of perception and prediction [11, 17, 47, 41], or adopts a direct optimization on planning without intermediate tasks [3, 4, 34], which lack interpretability and are hard to optimize. Recently, UniAD [11] presents a planning-oriented model which integrates various tasks in the dense BEV-Centric paradigm, achieving convincing performance. VAD [17] learns vectorized scene representations and improves planning safety with explicit constraints. GraphAD [47] constructs the interaction scene graph to model both dynamic and static relations. SparseDrive [36] introduces the symmetric sparse perception for parallel motion planner. However, using straightforward designs and exhaustive modeling without ego-centric interaction, will inevitably lead to unsatisfactory planning performance and inferior efficiency.
3 Our Approach
3.1 Overview
The overall framework of proposed DiFSD is illustrated in Fig. 2, which deals with the end-to-end planning task in an ego-centric fully sparse paradigm. Specifically, DiFSD mainly consists of four parts: visual encoder, sparse perception, hierarchical interaction and iterative motion planner. First, the visual encoder extracts multi-scale spatial features from given multi-view images. Then the sparse perception takes the encoded features as input to perform detection, tracking and online mapping simultaneously. In the hierarchical interaction module, the ego query equipped with a geometric prior is introduced to pick out the interactive queries through ego-centric cross attention and hierarchical selection. In the iterative motion planner, both interactive agents and ego-vehicle are considered for joint motion prediction, then the predicted multi-modal ego-trajectories are further optimized iteratively. Meanwhile, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are conducted for uncertainty modeling of motion and planning tasks respectively.
3.2 Problem Formulation
Given multi-view camera image sequence can be denoted as , where is the number of camera views and indicates the temporal length till current timestep respectively. Annotation of input for end-to-end planning is composed by a set of future waypoints of the ego-vehicle , where = 3 is the planning time horizon, and (, ) is the BEV location transformed to the ego-vehicle coordinate system at current timestep . Meanwhile, driving command as well as ego-status is also provided. Annotation set is used during training. During prediction, the planned trajectory of ego-vehicle should fit the annotation with minimum L2 errors and collision rate with surrounding agents.
3.3 Sparse Perception
After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [24, 25] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [36].
Detection. Following the previous sparse perception methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box is denoted as , which contains location, dimension, yaw angle as well as velocity respectively. Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of . The updated instance features are adopted to predict the classification scores and box offsets respectively. Temporal instance denoising is introduced to improve model stability.
Online Mapping. Similarly, we adopt an additional detection branch of same structure for online mapping. Differently, the geometric anchor of each static map element is denoted as points. Therefore, surrounding maps can be represented by a group of map instance features and anchor polylines .
3.4 Ego-Env Hierarchical Interaction
We continue to perform hierarchical interaction between the ego-vehicle and surrounding objects. As shown in Fig. 2, the hierarchical interaction module mainly consists of three parts: Ego-Object Dual Interaction, Intention-guided Geometric Attention and Coarse-to-Fine Selection.
Ego-Object Dual Interaction. As shown in Fig. 3, a learnable embedding is randomly initialized to serve as ego query, along with an ego anchor box together to represent the ego-vehicle. Both ego-centric cross attention with surrounding objects ( and object-centric self attention are conducted consecutively to capture the mutual information comprehensively. During the attention calculation process, we combine positional embedding and query feature in a concatenated manner instead of an additive approach, which can effectively retain both semantic and geometric clues for interaction modeling.
Intention-Guided Geometric Attention. To enhance the accuracy and explainability of query ranking to facilitate selection, we introduce an ego-centric geometric prior additionally. As shown in Fig. 2, the intention-guided attention module is adopted to assess the importance of surrounding agent and map queries, which mainly consists of three steps: Response Map Learning, Reference Line Generation and Interactive Score Fusion.
Specifically, we use four MLPs to encode the ego-intention respectively, including velocity, acceleration, angular velocity and driving command. And then we concatenate these embeddings to obtain ego-intention features , which are further concatenated with the position embeddings of a group of pre-defined locations to cover densely distributed grid cells in the BEV plane. The position of each grid cell is represented as . Finally, the concatenated geometric features are fed to a single SE [9] block to learn response map , which is supervised by the normalized minimum distance from to the ego future waypoints. The motivation is that the Closest In-Path Vehicle / Stationary are prone to affect the ego-intention, and vice versa.
With the predicted response map , we first generate the reference line through row-wise thresholding, which are further used to generate the normalized distance map (See Fig. 2). Then we can obtain the geometric score for each surrounding query by referring to the . The reason why we don’t get the geometric score from directly is that the imbalanced distribution of ego-intention and future waypoints may lead to the inferior quality of .
Finally, as shown in Fig. 4, we perform interactive score fusion through multiplying the attention, geometric and classification scores during the ego-centric cross attention:
where the distance-prior is weighted with the attention score for both interaction and selection. is inner product, is dot product, and is the channel dimension.
Coarse-to-Fine Selection. To capture the interaction information from coarse to fine, we stack dual-interaction layers in a cascaded manner, where a top-K operation is appended between each two consecutive layers, thus the interactive objects can be gradually selected for latter prediction and planning usages. We claim that only a few interactive objects need to be considered for motion prediction, which are enough yet efficient for ego-centric path planning, instead of all detected agents existing in the driving scene.
3.5 Iterative Motion Planner
As shown in Fig. 2, the iterative motion planner is designed to conduct motion prediction for both interactive agents and ego-vehicle, and then optimize the proposal ego-trajectory with both safety and kinematic constrains iteratively.
Joint Motion Prediction. With regard to the trajectory prediction, both surrounding agents and ego-vehicle are adopted for motion modeling in a joint decoder, unlike previous works [11, 41, 17] which neglect the crucial interactions between near agents and ego-vehicle when making motion predictions, especially in the common scenarios like intersections. Another difference is that only the interactive objects (CIPV) sparsely selected in the former module are considered, instead of all detected agents in the driving scene which maybe irrelevant to the ego-vehicle planning. As for the joint motion decoder, we prepare three copies of ego query to indicate different driving intentions (i.e., turn left, turn right and keep forward), which are combined with to conduct agent-level self attention and agent-map cross attention respectively. And then we concatenate these output attended features to predict multi-modal trajectories , and classification scores , for both agents and ego-vehicle, where is the number of driving command for planning, are the mode number, are the future timestamps.
Planning Optimization. With the predicted multi-intention and multi-modal trajectories of ego-vehicle, we can select the most probable proposal trajectory with the input driving command and classification score . As shown in Fig. 3(b), ego-agent, ego-map and ego-navigator cross attentions are conducted consecutively for planning optimization, where the offsets for each future waypoint are predicted upon the proposal trajectory respectively with several planning constraints proposed in [17] to ensure safety.
Iterative Refinement. To further promote the stability and performance of the whole end-to-end system, an additional iterative refinement strategy is proposed to continuously update the reference line and distance map with refined ego trajectory as illustrated in Fig. 2, thus ensuring the interaction quality and selection accuracy of interactive queries.
3.6 Uncertainty Denoising
Due to the planning-oriented modular design, output uncertainty from each individual module will be inevitably introduced and passed through to the downstream tasks, leading to inferior and fragile system. Under this circumstance, we propose a two-level uncertainty modeling strategy to further stabilize the whole framework. On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents equipped with groups of random noises following uniform distributions. locates within two different ranges of and to indicate positives and negatives respectively, where indicates the noise scale. This process aims to promote the stability of motion forecasting for interactive agents with uncertain detected positions, scales and velocities. On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage. Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
3.7 End-to-End Learning
Multi-stage Training. To facilitate the model convergence and training performance, we divide the training process into two stages. In stage-1, the sparse perception, hierarchical interaction and joint motion prediction tasks are trained from scratch to learn sparse scene representation, interaction and motion capability respectively. Note that no selection operation is adopted in stage-1, namely all detected agents are considered for motion forecasting to make full use of annotations. In stage-2, the geometric attention module and the iterative planning optimizer are added to train jointly for overall optimization with uncertainty modeling.
Loss Functions. The overall optimization function mainly includes five tasks, where each task can be optimized with both classification and regression losses. The overall loss function for end-to-end training can be formulated as:
where is a combination of binary classification loss and regression loss to learn geometric score, where the positive (interactive) samples are denoted as grid cells with geometric score (within 3 for each future waypoint). An additional regression loss is included in for ego status prediction, instead of directly using it as input to the planner as [11, 41, 17], which will lead to information leakage as proven in [21]. Meanwhile, vectorized planning constrains identified in [17] such as collision, overstepping and direction are also included in for regularization. is the number of motion planning stages.
4 Experiments
4.1 Datasets and Metrics
Our experiments are first conducted on the challenging public nuScenes [1] dataset, which contains 1000 driving scenes lasting 20 seconds respectively. Over 1.4M 3D bounding boxes of 23 categories are provided in total, which are annotated at 2Hz. Following the conventions [11, 17], Collision Rate () and L2 Displacement Error (DE) () are adopted to measure the open-loop planning performance. To study the effect of various perception encoders, we also evaluate the 3D object detection and online mapping results using mAP and NDS metrics respectively. Besides, Bench2Drive [15] provides a comprehensive benchmarking for evaluating multiple abilities of end-to-end AD systems in a closed-loop manner, which collects 1000 clips covering 44 interactive scenarios, 23 weathers and 12 towns in CARLA v2 [5]. Following the official settings, we use 950 clips for training while leaving 50 clips for open-loop evaluation. As for the closed-loop evaluation, we run the trained model in CARLA with 220 test routes and calculate the closed-loop metrics such as Driving Score (DS), Success Rate (SR) and Efficiency, respectively.
4.2 Implementation Details
DiFSD plans a 3 future ego-trajectory with 2 history information as input. Our DiFSD has two variants, namely DiFSD-S and DiFSD-B. As for DiFSD-S, both perspective-view perception version DiFSD-S (PV) and BEV perception version DiFSD-S (BEV) are all implemented for comparison. ResNet50 [8] is adopted as the default backbone network for visual encoding. The perception range is set to 6030 longitudinally and laterally. Input image size of DiFSD-S is resized to . For DiFSD-S (BEV), the default number of BEV query, map query, agent query is , and 300, respectively. For DiFSD-S (PV), is 6, is 900 and is 100 respectively. Each map element contains 20 map points. The feature dimension is 256. The noise scale is set to 2.0 and 0.2FD for motion and planning respectively. is set to 3. DiFSD-B has larger input image resolution ( and backbone network (ResNet101). We use AdamW [30] optimizer and Cosine Annealing [29] scheduler to train DiFSD with weight decay 0.01 and initial learning rate -4. DiFSD is trained for 60 epochs and 4 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A100 GPUs with batch size 8 empirically.
| Protocol | Method | Backbone | L2 (m) | Collision (%) | Latency (ms) | FPS | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
| ST-P3 Metrics | ST-P3 [10] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 |
| OccNet [37] | ResNet101-DCN | 1.29 | 2.13 | 2.99 | 2.13 | 0.21 | 0.59 | 1.37 | 0.72 | - | - | |
| FusionAD∗ [41] | ResNet101+SECOND [39] | - | - | - | 1.03 | 0.25 | 0.13 | 0.25 | 0.21 | - | - | |
| VAD-Base [17] | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.07 | 0.17 | 0.41 | 0.22 | 224.3 | 4.5 | |
| SparseDrive-S [36]† | ResNet50 | 0.30 | 0.58 | 0.95 | 0.61 | 0.47 | 0.47 | 0.69 | 0.54 | 111.1 | 9.0 | |
| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 | 67.7 | 14.8 | |
| SparseDrive Metrics‡ | UniAD [11]† | ResNet101-DCN | 0.45 | 0.70 | 1.04 | 0.73 | 0.62 | 0.58 | 0.63 | 0.61 | 555.6 | 1.8 |
| VAD-Base [17]† | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.03 | 0.19 | 0.43 | 0.21 | 224.3 | 4.5 | |
| SparseDrive-S [36] | ResNet50 | 0.29 | 0.58 | 0.96 | 0.61 | 0.01 | 0.05 | 0.18 | 0.08 | 111.1 | 9.0 | |
| SparseDrive-B [36] | ResNet101 | 0.29 | 0.55 | 0.91 | 0.58 | 0.01 | 0.02 | 0.13 | 0.06 | 137.0 | 7.3 | |
| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
| DiFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 93.7 | 10.7 | |
| DiFSD-B (PV) | ResNet101 | 0.15 | 0.30 | 0.54 | 0.32 | 0.00 | 0.04 | 0.15 | 0.06 | 119.6 | 8.4 | |
4.3 Main Results
Open-loop Planning Evaluation. As show in Tab. 1, DiFSD shows great advantages in both performance and efficiency compared with previous works. On one hand, DiFSD-S achieves the minimum L2 error even with lightweight visual backbone and inferior BEV perception encoder. Specifically, compared with BEVFormer-based end-to-end methods [11, 17], DiFSD-S (BEV) reduces the average L2 error by a great margin (0.38 and 0.37, separately), while significantly reducing the average collision rates by 84% and 52% respectively. Equipped with deeper visual backbone and advanced sparse detectors from Perspective View (PV), the average L2 error and collision rates can be further reduced to 0.32 and to 0.06 respectively. Notably, we are the first to achieve 0 collision rate on 1. On the other hand, benefiting from the ego-centric hierarchical interaction, only sparse interactive agents (2%) are considered for motion planning. Hence, DiFSD-S can achieve great efficiency with 14.8 FPS, and faster than UniAD [11] and VAD [17] respectively.
Closed-loop Planning Evaluation. We further validate the closed-loop performance in Bench2Drive [15], which has been proposed recently for comprehensive benchmarking of end-to-end planning methods. As shown in Tab. 2, AD-MLP [45] has a high L2 error and bad closed-loop planning performance using merely ego status as input, which is different from findings in nuScenes [1], demonstrating the behavior diversity in Bench2Drive. UniAD [11] has a lower L2 error compared to VAD [17] but with worse closed-loop planning performance as discussed in [21]. Notably, DiFSD-S achieves both the lowest L2 error and best closed-loop performance with great efficiency, showcasing the superiority and generalizability of our proposed method.
4.4 Ablation Study
We conduct extensive experiments to study the effectiveness of our DiFSD. We use DiFSD-S as default for ablation. More experiments are provided in the Appendix.
Necessity of Geometric Prior. We claim that the Closest In-Path Vehicle as well as Stationary (CIPV / CIPS) are more likely to interact with the ego-vehicle. To verify the necessity of such geometric prior, we conduct exhaustive ablations of the ego-centric query selector as show in Tab. 3. Without ego-centric selection, fewer objects randomly selected can result in worse planning results. While using the ego-centric cross attention, only 2% of surrounding queries are enough for achieving convincing planning performance. Besides, introducing the geometric prior through attention can dramatically reduce the L2 error and collision rate by 8% and 42% respectively. Meanwhile, when utilizing the ground-truth geometric score for upper-limit evaluation, we can obtain the extremely excellent planning performance (0.23 average L2 error). Undoubtedly, the proposed ego-centric query selector equipped with geometric attention is nontrivial for motion planner.
Effect of designs in Hierarchical Interaction. Tab. 4 shows the effectiveness of our elaborate designs in the hierarchical interaction module, which contains three main designs such as Dual Interaction (DI), Geometric Attention (GA) and Coarse-to-Fine Selection (CFS). DI models both ego-centric and object-centric interactions respectively, which improves the planning performance greatly as expected. GA facilitates the query selection process as discussed in Tab. 3, which reduces the collision rate by a great margin (42%). And CFS contributes to the interaction modeling quality through hierarchical receptive fields from global to local. All of these three designs combined together can achieve overall convincing planning performance.
| Object Selection | Geometric Attention | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | ||
| 100% | ✗ | 0.27 | 0.47 | 0.74 | 0.49 | 0.10 | 0.21 | 0.37 | 0.22 |
| Random (5%) | ✗ | 0.28 | 0.49 | 0.79 | 0.52 | 0.08 | 0.17 | 0.38 | 0.21 |
| Random (2%) | ✗ | 0.33 | 0.57 | 0.87 | 0.59 | 0.18 | 0.30 | 0.51 | 0.33 |
| 0% | ✗ | 2.25 | 3.75 | 5.26 | 3.75 | 2.82 | 5.42 | 6.39 | 4.88 |
| Attn (5%) | ✗ | 0.16 | 0.34 | 0.63 | 0.38 | 0.07 | 0.09 | 0.31 | 0.16 |
| Attn (2%) | ✗ | 0.16 | 0.34 | 0.61 | 0.37 | 0.08 | 0.11 | 0.27 | 0.15 |
| Attn (2%) | Random | 0.17 | 0.36 | 0.67 | 0.40 | 0.07 | 0.10 | 0.34 | 0.17 |
| Attn (2%) | GroundTruth | 0.14 | 0.23 | 0.33 | 0.23 | 0.07 | 0.08 | 0.10 | 0.07 |
| Attn (2%) | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
| DI | GA | CFS | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||
| ✗ | ✓ | ✓ | 0.18 | 0.35 | 0.62 | 0.38 | 0.09 | 0.12 | 0.23 | 0.14 |
| ✓ | ✗ | ✓ | 0.16 | 0.34 | 0.61 | 0.37 | 0.08 | 0.11 | 0.27 | 0.15 |
| ✓ | ✓ | ✗ | 0.16 | 0.33 | 0.59 | 0.36 | 0.09 | 0.11 | 0.25 | 0.15 |
| ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
Effect of designs in Motion Planner. As for motion planner in DiFSD, Joint Motion Prediction (JMP), Planning Optimization (PO) as well as Iterative Refinement (IR) makes up the planning pipeline of ego-vehicle. Besides, Uncertain Denoising (UD) contributes to the system stability and training convergence. Tab. 5 explores the effect of each design exhaustively. ID-1 indicates evaluating the proposal trajectory of ego-vehicle predicted together with interactive agents, which achieves competitive L2 error but is easier to collide with surrounding agents. ID-2 improves the collision rate greatly by 38.9% with the help of PO and planning constraints [17] during training phase. ID-4 emphasizes the importance of IR in improving the quality of ego-planning trajectory (average 5.4% L2 error and 36.3% collision rate reduction respectively). ID-3 reflects the benefit of UD used for end-to-end training compared to ID-4.
Effect of Iterative Refinement stages. We continue to study the number of refinement stages in Tab. 6. We can observe that our DiFSD can obtain superior planning performance with one additional refinement stage (36.3% collision rate reduction), which becomes saturated when introducing more stages. Hence, two-stage interacted motion planner is enough for achieving convincing results.
Effect of Uncertainty Denoising. We also validate the effectiveness of uncertainty denoising strategy including position-level motion diffusion and trajectory-level planning denoising. As shown in Tab. 7, motion diffusion can improve the prediction stability with uncertain agent positions, while the planning denoising can also strengthen the trajectory regression precision of ego-vehicle.
| ID | JMP | PO | IR | UD | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
| 1 | ✓ | ✗ | ✗ | ✓ | 0.23 | 0.48 | 0.83 | 0.51 | 0.08 | 0.13 | 0.35 | 0.18 |
| 2 | ✓ | ✓ | ✗ | ✓ | 0.16 | 0.33 | 0.61 | 0.37 | 0.01 | 0.08 | 0.23 | 0.11 |
| 3 | ✓ | ✓ | ✓ | ✗ | 0.16 | 0.34 | 0.64 | 0.38 | 0.07 | 0.07 | 0.17 | 0.10 |
| 4 | ✓ | ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
| Number of stages | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |
| 1 | 0.16 | 0.33 | 0.61 | 0.37 | 0.01 | 0.08 | 0.23 | 0.11 |
| 2 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
| 3 | 0.16 | 0.33 | 0.60 | 0.36 | 0.01 | 0.40 | 0.22 | 0.09 |
| 4 | 0.16 | 0.33 | 0.61 | 0.36 | 0.00 | 0.04 | 0.20 | 0.08 |
Runtime of each module. We evaluate the modular runtime of DiFSD-S with different perception paradigms as shown in Tab. 8. Generally, visual backbone and perception encoder occupy the most of the runtime for feature extraction and scene understanding. Hierarchical interaction also takes a significant part for interaction modeling and interactive query selection. Thanks to the ego-centric interaction and selection module, the joint motion planner only consumes 7.9ms to plan the future ego-trajectory. Notably, though PV-based perception module [24] achieves superior detection performance, it consumes more computation cost due to the repeated query projection, deformable feature aggregation as well as the symmetric sparse perception architecture for unified object detection and online mapping, without a shared BEV feature as [17, 11].
4.5 Qualitative Results
We visualize the motion trajectories of sparse interactive agents as well as planning results of DiFSD as illustrated in Fig. 5. Both surrounding camera images and prediction results on BEV are provided accordingly. Besides, we also project the planning trajectories to the front-view camera image. Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of DiFSD motivation. DiFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
| Position Diffusion | Trajectory Denoising | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | ||
| ✗ | ✗ | 0.16 | 0.34 | 0.64 | 0.38 | 0.07 | 0.07 | 0.17 | 0.10 |
| ✓ | ✗ | 0.16 | 0.34 | 0.63 | 0.37 | 0.02 | 0.04 | 0.15 | 0.07 |
| ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
| Module | Bird’s-Eye View (BEV) Perception | Perspective-View (PV) Perception | ||
| Latency (ms) | Proportion (%) | Latency (ms) | Proportion (%) | |
| Backbone | 8.4 | 12.4 | 11.0 | 11.8 |
| Perception Encoder | 34.4 | 50.8 | 54.9 | 58.6 |
| Hierarchical Interaction | 17.0 | 25.1 | 19.9 | 21.2 |
| Motion Prediction | 4.5 | 6.6 | 4.5 | 4.8 |
| Planning Optimization | 3.4 | 5.1 | 3.4 | 3.6 |
| Total | 67.7 | 100 | 93.7 | 100 |
5 Conclusion
In this paper, we propose a fully sparse paradigm for end-to-end self-driving in an ego-centric manner, termed as DiFSD. DiFSD revisits the human driving behavior and conducts hierarchical interaction based on sparse representation and perception results. Only interactive agents are considered for joint motion prediction with the ego-vehicle. Iterative planning optimization strategy contributes to the driving safety with high efficiency. Besides, uncertainty modeling is conducted to improve the stability of end-to-end system. Extensive ablations and comparisons reveal the superiority and great potential of our ego-centric fully sparse paradigm for future research.
Supplementary Material
Appendix A Evaluation Metrics
Perception. The evaluation for detection and tracking follows standard evaluation protocols [1]. For detection, we use mean Average Precision(mAP), mean Average Error of Translation(mATE), Scale(mASE), Orientation(mAOE), Velocity(mAVE), Attribute(mAAE) and nuScenes Detection Score(NDS) to evaluate the model performance. For online mapping, we calculate the Average Precision(AP) of three map classes: lane divider, pedestrian crossing and road boundary, then average across all classes to get mean Average Precision (mAP).
Planning. We adopt commonly used L2 error and collision rate to evaluate the planning performance. The evaluation of L2 error is aligned with VAD [17]. For collision rate, there are two drawbacks in previous [11, 17] implementation, resulting in inaccurate evaluation in planning performance. On one hand, previous benchmark convert obstacle bounding boxes into occupancy map with a grid size of 0.5m, resulting in false collisions in certain cases, e.g. ego vehicle approaches obstacles that smaller than a single occupancy map pixel [45]. (2) The heading of ego vehicle is not considered and assumed to remain unchanged [21]. To accurately evaluate the planning performance, we account for the changes in ego heading by estimating the yaw angle through trajectory points, and assess the presence of a collision by examining the overlap between the bounding boxes of ego vehicle and obstacles. We reproduce the planning results on our benchmark with official checkpoints [11, 17] for a fair comparison.
Appendix B More Ablation Study
Necessity and Order of Object Selection. Tab. 1 studies the necessity of agent and map selection during the ego-centric hierarchical interaction. We can observe that agent selection contributes more than the map selection, especially in the driving safety. And both of agent and map interactions are conducted in a cascaded order is inferior than the parallel manner, where the updated ego query from parallel outputs are concatenated for joint motion prediction.
| Agent Selection | Map Selection | Cascade | Parallel | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | ||||
| ✓ | ✗ | - | - | 0.16 | 0.34 | 0.64 | 0.38 | 0.03 | 0.05 | 0.22 | 0.10 |
| ✗ | ✓ | - | - | 0.17 | 0.35 | 0.63 | 0.38 | 0.02 | 0.06 | 0.28 | 0.12 |
| ✓ | ✓ | ✓ | - | 0.16 | 0.34 | 0.62 | 0.37 | 0.05 | 0.07 | 0.30 | 0.14 |
| ✓ | ✓ | - | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
Effect of Interactive Score Fusion. During the ego-centric query selection, both geometric and classification scores are considered to ensure that the selected closest in-path queries are true positive agents or maps, which are adopted for motion planner. Tab. 2 shows the effect of three types of scores used for query ranking, namely attention, geometry and confidence scores. As Eq. 1, interactive score obtained by multiplying these three scores can achieve the best selection quality and planning performance. without confidence score fails to distinguish between background and foreground queries, resulting in inferior performance.
| Attention Score | Geometric Score | Classification Score | Planning L2 (m) | Planning Coll. (%) | ||||||
| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||
| ✓ | ✗ | ✗ | 0.18 | 0.36 | 0.66 | 0.39 | 0.09 | 0.11 | 0.28 | 0.16 |
| ✓ | ✓ | ✗ | 0.17 | 0.35 | 0.65 | 0.38 | 0.01 | 0.07 | 0.24 | 0.11 |
| ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
Effect of Sparse Perception. Recent end-to-end planning method [36] resorts to the sparse perception fashion to provide advanced 3D detection and online mapping results with high efficiency. To study the significance of advanced perception encoders for ego-planning, we compare the perception performance of various end-to-end methods as shown in Tab. 3. With sparse perception encoder [24], the performance of 3D object detection and online mapping can be greatly improved (10.6 NDS and 7.5 mAP, respectively) compared with dense BEV-based perception paradigm [20]. And the end-to-end planner [36] equipped with the advanced perception encoder can consistently boost the planning performance. Therefore, the perception performance is essential for the end-to-end planner.
| Method | Backbone | BEV | mAP | NDS |
| BEVFormer [20] | ResNet101-DCN | ✓ | 41.6 | 51.7 |
| Sparse4Dv3 [24] | ResNet101∗ | ✗ | 53.7 | 62.3 |
| UniAD [11] | ResNet101-DCN | ✓ | 38.0 | 49.8 |
| VAD† [17] | ResNet50 | ✓ | 27.3 | 39.7 |
| SparseDrive-S [36] | ResNet50 | ✗ | 41.8 | 52.5 |
| SparseDrive-B [36] | ResNet101* | ✗ | 49.6 | 58.8 |
| DiFSD-S | ResNet50 | ✓ | 32.8 | 45.8 |
| DiFSD-S | ResNet50 | ✗ | 41.0 | 52.8 |
| DiFSD-B | ResNet101* | ✗ | 49.6 | 58.9 |
| Method | APped | APdivider | APboundary | mAP |
| VectorMapNet [28] | 36.1 | 47.3 | 39.3 | 40.9 |
| MapTR [23] | 56.2 | 59.8 | 60.1 | 58.7 |
| VAD† [17] | 40.6 | 51.5 | 50.6 | 47.6 |
| SparseDrive-S [36] | 49.9 | 57.0 | 58.4 | 55.1 |
| SparseDrive-B [36] | 53.2 | 56.3 | 59.1 | 56.2 |
| DiFSD-S (BEV) | 46.7 | 54.3 | 56.0 | 52.3 |
| DiFSD-S (PV) | 54.9 | 55.7 | 57.3 | 56.0 |
| DiFSD-B (PV) | 52.3 | 58.2 | 59.3 | 56.6 |
Appendix C Analysis & Discussion
The GroundTruth future state distribution of ego-vehicle on nuScenes validation set is illustrated in Fig. A1, which is calculated with fixed time interval (1) between consecutive predicted waypoints. And we also compare the output ego-state distribution of different popular end-to-end methods based on planned trajectories respectively, as show in Fig. A2. We can observe that without ego-centric design, the optimized end-to-end model is unable to handle various emergencies appearing in the driving scenarios, where the absolute values of and are larger than normal situations. Under this circumstance, the output planned trajectories cannot conform to the expert routes as expected. However, our DiFSD performs consistently better in planning the future ego states with variable speed and acceleration, owing to the ego-centric hierarchical interaction and selection mechanism, thus the iterative motion planner can focus on the interactive agents rather than irrelevant objects.
Appendix D Visualization
As show in Fig. A3-A5, we provide more visualization results to illustrate the effectiveness of DiFSD on various driving scenarios (i.e., interactive / non-interactive scenes, lane-change / lane-keep / following scenes, overtaking / avoidance scenes, intersection scenes) under different commands (i.e., “Go Straight”, “Turn Left”, “Turn Right”).
Moreover, we also observe some failure cases on nuScenes validation set as illustrated in Fig. A6 and Fig. A7. The corresponding explanations as well as analysis are described in the captions below respectively, which mainly attribute to the irrationality of the expert (ground truth) future trajectories of ego vehicle, demonstrating the strong generalizability of our proposed DiFSD in planning the efficient and reasonable results based on the selected sparse representations of surrounding driving scenarios.

### v4

@@
 As for the sparse fashion, current methods [36, 26, 27, 25, 24] adopt a set of sparse queries to integrate spatial-temporal aggregations from multi-view image feature sequence for iterative anchor refinement.
-Besides, Multi-Object Tracking (MOT) across multi-cameras is also required for downstream tasks.
-Recent works [44, 46, 42, 32, 35] seek to explore the joint detection and tracking methods by introducing track queries to detect the unique instances continuously.
-Sparse4Dv3 [24] proposes an advanced 3D detector which takes the spatial-temporal information to propagate temporal instances for identity reserving, thus achieving superior end-to-end performance without additional tracking designs.
 2.2 Online Mapping
@@
 After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [24, 25] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [34].
-Detection.
-Following the previous sparse perception methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively.
-And each anchor box is denoted as , which contains location, dimension, yaw angle as well as velocity respectively.
+Detection. Following the previous methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted as:
+which contains location, dimension, yaw angle as well as velocity respectively.
 Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of .
@@
 Under this circumstance, we propose a two-level uncertainty modeling strategy to further stabilize the whole framework.
-On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents equipped with groups of random noises following uniform distributions.
+On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents:
+which are equipped with groups of random noises following uniform distributions.
 locates within two different ranges of and to indicate positives and negatives respectively, where indicates the noise scale.
@@
 Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
+| Protocol | Method | Backbone | L2 (m) | Collision (%) | Latency (ms) | FPS | ||||||
+| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
+| ST-P3 Metrics | ST-P3 [10] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 |
+| OccNet [35] | ResNet101-DCN | 1.29 | 2.13 | 2.99 | 2.13 | 0.21 | 0.59 | 1.37 | 0.72 | - | - | |
+| FusionAD∗ [39] | ResNet101+SECOND [37] | - | - | - | 1.03 | 0.25 | 0.13 | 0.25 | 0.21 | - | - | |
+| VAD-Base [17] | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.07 | 0.17 | 0.41 | 0.22 | 224.3 | 4.5 | |
+| SparseDrive-S [34]† | ResNet50 | 0.30 | 0.58 | 0.95 | 0.61 | 0.47 | 0.47 | 0.69 | 0.54 | 111.1 | 9.0 | |
+| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 | 67.7 | 14.8 | |
+| SparseDrive Metrics‡ | UniAD [11]† | ResNet101-DCN | 0.45 | 0.70 | 1.04 | 0.73 | 0.62 | 0.58 | 0.63 | 0.61 | 555.6 | 1.8 |
+| VAD-Base [17]† | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.03 | 0.19 | 0.43 | 0.21 | 224.3 | 4.5 | |
+| SparseDrive-S [34] | ResNet50 | 0.29 | 0.58 | 0.96 | 0.61 | 0.01 | 0.05 | 0.18 | 0.08 | 111.1 | 9.0 | |
+| SparseDrive-B [34] | ResNet101 | 0.29 | 0.55 | 0.91 | 0.58 | 0.01 | 0.02 | 0.13 | 0.06 | 137.0 | 7.3 | |
+| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
+| DiFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 93.7 | 10.7 | |
+| DiFSD-B (PV) | ResNet101 | 0.15 | 0.30 | 0.54 | 0.32 | 0.00 | 0.04 | 0.15 | 0.06 | 119.6 | 8.4 | |
 3.7 End-to-End Learning
@@
 DiFSD is trained for 60 epochs and 4 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A100 GPUs with batch size 8 empirically.
-| Protocol | Method | Backbone | L2 (m) | Collision (%) | Latency (ms) | FPS | ||||||
-| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
-| ST-P3 Metrics | ST-P3 [10] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 |
-| OccNet [37] | ResNet101-DCN | 1.29 | 2.13 | 2.99 | 2.13 | 0.21 | 0.59 | 1.37 | 0.72 | - | - | |
-| FusionAD∗ [41] | ResNet101+SECOND [39] | - | - | - | 1.03 | 0.25 | 0.13 | 0.25 | 0.21 | - | - | |
-| VAD-Base [17] | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.07 | 0.17 | 0.41 | 0.22 | 224.3 | 4.5 | |
-| SparseDrive-S [36]† | ResNet50 | 0.30 | 0.58 | 0.95 | 0.61 | 0.47 | 0.47 | 0.69 | 0.54 | 111.1 | 9.0 | |
-| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 | 67.7 | 14.8 | |
-| SparseDrive Metrics‡ | UniAD [11]† | ResNet101-DCN | 0.45 | 0.70 | 1.04 | 0.73 | 0.62 | 0.58 | 0.63 | 0.61 | 555.6 | 1.8 |
-| VAD-Base [17]† | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.03 | 0.19 | 0.43 | 0.21 | 224.3 | 4.5 | |
-| SparseDrive-S [36] | ResNet50 | 0.29 | 0.58 | 0.96 | 0.61 | 0.01 | 0.05 | 0.18 | 0.08 | 111.1 | 9.0 | |
-| SparseDrive-B [36] | ResNet101 | 0.29 | 0.55 | 0.91 | 0.58 | 0.01 | 0.02 | 0.13 | 0.06 | 137.0 | 7.3 | |
-| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
-| DiFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 93.7 | 10.7 | |
-| DiFSD-B (PV) | ResNet101 | 0.15 | 0.30 | 0.54 | 0.32 | 0.00 | 0.04 | 0.15 | 0.06 | 119.6 | 8.4 | |
 4.3 Main Results
@@
 As Eq.
-1, interactive score obtained by multiplying these three scores can achieve the best selection quality and planning performance.
+2, interactive score obtained by multiplying these three scores can achieve the best selection quality and planning performance.
 without confidence score fails to distinguish between background and foreground queries, resulting in inferior performance.

### v5

@@
-DiFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving
+EgoFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving
 Abstract
-Current end-to-end autonomous driving methods resort to unifying modular designs for various tasks (e.g.
+Current End-to-End Autonomous Driving (E2E-AD) methods resort to unifying modular designs for various tasks (e.g.
 perception, prediction and planning).
-Although optimized in a planning-oriented spirit with a fully differentiable framework, existing end-to-end driving systems without ego-centric designs still suffer from unsatisfactory performance and inferior efficiency, owing to the rasterized scene representation learning and redundant information transmission.
-In this paper, we revisit the human driving behavior and propose an ego-centric fully sparse paradigm, named DiFSD, for end-to-end self-driving.
-Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner.
-The sparse perception module performs detection, tracking and online mapping based on sparse representation of the driving scene.
+Although optimized with a fully differentiable framework in a planning-oriented manner, existing end-to-end driving systems lacking ego-centric designs still suffer from unsatisfactory performance and inferior efficiency, due to rasterized scene representation learning and redundant information transmission.
+In this paper, we propose an ego-centric fully sparse paradigm, named EgoFSD, for end-to-end self-driving.
+Specifically, EgoFSD consists of sparse perception, hierarchical interaction and iterative motion planner.
+The sparse perception module performs detection and online mapping based on sparse representation of the driving scene.
 The hierarchical interaction module aims to select the Closest In-Path Vehicle / Stationary (CIPV / CIPS) from coarse to fine, benefiting from an additional geometric prior.
 As for the iterative motion planner, both selected interactive agents and ego-vehicle are considered for joint motion prediction, where the output multi-modal ego-trajectories are optimized in an iterative fashion.
-Besides, both position-level motion diffusion and trajectory-level planning denoising are introduced for uncertainty modeling, thus facilitating the training stability and convergence of the whole framework.
-Extensive experiments conducted on nuScenes and Bench2Drive datasets demonstrate the superior planning performance and great efficiency of DiFSD.
+In addition, position-level motion diffusion and trajectory-level planning denoising are introduced for uncertainty modeling, thereby enhancing the training stability and convergence speed.
+Extensive experiments are conducted on nuScenes and Bench2Drive datasets, which significantly reduces the average L2 error by 59% and collision rate by 92% than UniAD while achieves 6.9 faster running efficiency.
 1 Introduction
@@
 Recently, end-to-end paradigm integrates all tasks into a unified model for planning-oriented optimization, showcasing great potential in pushing the limit of autonomous driving performance.
-Literally, existing end-to-end models [11, 39, 17, 34] designed for reliable trajectory planning can be classified into two mainstreams as summarized in Fig.
+Literally, existing end-to-end models [12, 47, 17, 39, 44, 45, 46] designed for reliable trajectory planning can be classified into two mainstreams as summarized in Fig.
 1(a) and (b).
@@
 Therefore, the planning performance remains unsatisfactory in both planning safety, comfort and personification.
-To this end, we propose DiFSD, an Ego-Centric fully sparse paradigm as shown in Fig.
+To this end, we propose EgoFSD, an Ego-Centric fully sparse paradigm as shown in Fig.
 1(c).
-Specifically, DiFSD mainly consists of sparse perception, hierarchical interaction and iterative motion planner.
-In the sparse perception module, multi-scale image features extracted from visual encoder are adopted for object detection, tracking and online mapping simultaneously in a sparse manner.
+Specifically, EgoFSD consists of sparse perception, hierarchical interaction and iterative motion planner.
+In the sparse perception module, multi-scale image features extracted from visual encoder are adopted for object detection and online mapping simultaneously in a sparse manner.
 Then the hierarchical interaction performs ego-centric and object-centric dual interaction to select the CIPV / CIPS with the help of an additional geometric prior.
@@
 It can not only model the uncertain positions of interactive agents for motion prediction, but also enhance the quality of trajectory refinement with arbitrary offsets.
-With above elaborate designs, DiFSD exhibits the great potential of fully sparse paradigm for end-to-end autonomous driving, which significantly reduces the average L2 error by 56% and collision rate by 92% than UniAD [11] respectively.
-Notably, our DiFSD-S achieves faster running efficiency as well.
+With above elaborate designs, EgoFSD exhibits the great potential of fully sparse paradigm for end-to-end autonomous driving, which significantly reduces the average L2 error by 59% and collision rate by 92% than UniAD [12] respectively.
+Notably, our EgoFSD-B achieves faster running efficiency as well.
 In sum, the main contributions of our work are as follows:
-We propose an ego-centric Fully Sparse paradigm for end-to-end self-Driving, named as DiFSD, without any computationally intensive dense scene representation learning and redundant environmental modeling, which is proven to be effective and efficient for ego-planning.
+We propose an Ego-Centric Fully Sparse paradigm for end-to-end self-Driving, named EgoFSD, without dense representation learning and redundant environmental modeling, which is proven to be effective for ego-planning.
 We introduce a geometric prior through intention-guided attention, where the Closest In-Path Vehicle / Stationary (CIPV / CIPS) are gradually picked out through ego-centric cross attention and selection.
-Besides, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are adopted for uncertainty modeling of motion planning respectively.
+Besides, both position-level diffusion of interactive agents and trajectory-level denoising of ego-vehicle are adopted for uncertainty motion modeling.
 2 Related Work
-2.1 End-to-End Perception
+Object Detection.
 Recent years witness remarkable progress achieved in multi-view 3D detection, which mainly build elaborate designs upon the dense BEV (Bird’s Eye View) or sparse query features.
 To generate BEV features, LSS [36] lifts 2D image features to 3D space using depth estimation results, which are then splatted into BEV plane.
-Follow-up works apply such operation to perform view transform for 3D detection task [14, 13, 19, 7].
+Follow-up works apply such operation to perform view transform for 3D detection task [14, 21, 7].
 Differently, some works [22, 43, 13] project a series of predefined BEV queries in 3D space to the image domain for feature sampling.
 As for the sparse fashion, current methods [42, 29, 30, 28, 27] adopt a set of sparse queries to integrate spatial-temporal aggregations from multi-view image feature sequence for iterative anchor refinement.
-2.2 Online Mapping
+Online Mapping.
 Maps could provide important static scenario information to ensure driving safety.
@@
 MapTR [25] regards map elements as a set of points with equivalent permutations, while StreamMapNet [48] adopts a temporal fusion strategy to enhance the performance.
-However, all of them reply on dense BEV features for online map construction, which is computationally intensive and not extensible to the sparse manner.
-2.3 End-to-End Motion Prediction
-Motion prediction of surrounding agents in an end-to-end fashion can relieve the accumulative error between standalone models.
-FaF [31] predicts both current and future bounding boxes from images using a single convolution network.
+However, all of them rely on dense BEV features for online map construction, which is computationally intensive and thus inefficient.
+Motion Prediction.
+Predicting agent future trajectories is essential for the autonomous vehicle to understand motion intention of surrounding agents.
+FaF [34] predicts both current and future boxes from images using a single deep network.
 IntentNet [2] attempts to reason high-level behavior and long-term trajectories simultaneously.
@@
 PIP [16] further proposes to replace HD-Map with local vectorized map.
-2.4 End-to-End Planning
-End-to-end planning paradigm either unites modules of perception and prediction [11, 17, 42, 39], or adopts a direct optimization on planning without intermediate tasks [3, 4, 33], which lack interpretability and are hard to optimize.
+End-to-End Planning.
+End-to-end planning paradigm either unites modules of perception and prediction [17, 51, 47, 35, 37], or adopts a direct optimization on planning without intermediate tasks [3, 4, 19], which lack interpretability and are hard to optimize.
 Recently, UniAD [12] presents a planning-oriented model which integrates various tasks in the dense BEV-Centric paradigm, achieving convincing performance.
@@
 GraphAD [51] constructs the interaction scene graph to model both dynamic and static relations.
-SparseDrive [34] introduces the symmetric sparse perception for parallel motion planner.
-However, using straightforward designs and exhaustive modeling without ego-centric interaction, will inevitably lead to unsatisfactory planning performance and inferior efficiency.
+SparseDrive [39] introduces the symmetric sparse perception for parallel motion planner, which consumes more computation cost due to the repeated query projection and deformable feature aggregation, without a shared 3D feature [12, 17].
+Besides, using straightforward designs and exhaustive modeling without ego-centric interaction, will inevitably lead to unsatisfactory planning performance and inferior efficiency.
 3 Our Approach
 3.1 Overview
-The overall framework of proposed DiFSD is illustrated in Fig.
+The overall framework of proposed EgoFSD is illustrated in Fig.
 2, which deals with the end-to-end planning task in an ego-centric fully sparse paradigm.
-Specifically, DiFSD mainly consists of four parts: visual encoder, sparse perception, hierarchical interaction and iterative motion planner.
+Specifically, EgoFSD mainly consists of four parts: visual encoder, sparse perception, hierarchical interaction and iterative motion planner.
 First, the visual encoder extracts multi-scale spatial features from given multi-view images.
@@
 3.3 Sparse Perception
-After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [24, 25] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [34].
-Detection. Following the previous methods [25, 24], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted as:
+After extracting the multi-view visual features from sensor images using the visual encoder [8], sparse query-based perception method proposed in [27, 29, 40] is extended to perform unified detection and online mapping in parallel with the symmetric architecture as adopted in [39].
+Detection. Following [27, 28], surrounding agents can be represented by a group of instance features and anchor boxes respectively. And each anchor box can be denoted:
 which contains location, dimension, yaw angle as well as velocity respectively.
-Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through deformable aggregation of sample features projected from key points of .
+Taking , and the visual features as input, decoders are adopted to consecutively refine the anchor boxes and update the instance features through iterative decoding.
 The updated instance features are adopted to predict the classification scores and box offsets respectively.
-Temporal instance denoising is introduced to improve model stability.
+Refer to [40], hybrid attention between current propagated queries and history memory queue is conducted for temporal modeling.
+Besides, temporal instance denoising is introduced to improve model stability.
 Online Mapping.
@@
 Specifically, we use four MLPs to encode the ego-intention respectively, including velocity, acceleration, angular velocity and driving command.
-And then we concatenate these embeddings to obtain ego-intention features , which are further concatenated with the position embeddings of a group of pre-defined locations to cover densely distributed grid cells in the BEV plane.
-The position of each grid cell is represented as .
-Finally, the concatenated geometric features are fed to a single SE [9] block to learn response map , which is supervised by the normalized minimum distance from to the ego future waypoints.
-The motivation is that the Closest In-Path Vehicle / Stationary are prone to affect the ego-intention, and vice versa.
-With the predicted response map , we first generate the reference line through row-wise thresholding, which are further used to generate the normalized distance map (See Fig.
+Then we concatenate these embeddings to obtain ego-intention features , which are further concatenated with the position embeddings of a group of pre-defined locations to cover densely distributed grid cells in the BEV plane.
+Finally, the concatenated geometric features are fed to a single SE [9] block to learn response map , which is supervised by the normalized minimum distance from to interpolated future trajectory with waypoints.
+Thus, for each grid , the response value can be formulated as:
+where is set to 30 for trajectory interpolation empirically. The motivation is that the Closest In-Path Vehicle / Stationary are prone to affect the ego-intention, and vice versa.
+With the predicted , we first generate the reference line through row-wise thresholding, which are further used to generate the normalized distance map (See Fig.
 2).
-Then we can obtain the geometric score for each surrounding query by referring to the .
-The reason why we don’t get the geometric score from directly is that the imbalanced distribution of ego-intention and future waypoints may lead to the inferior quality of .
+Then we can obtain the geometric score for each object query by referring to the .
+The reason why we don’t get the geometric score from directly is that the imbalanced distribution of ego-intention and future waypoints [23] may lead to the inferior quality of .
 Finally, as shown in Fig. 4, we perform interactive score fusion through multiplying the attention, geometric and classification scores during the ego-centric cross attention:
-where the distance-prior is weighted with the attention score for both interaction and selection. is inner product, is dot product, and is the channel dimension.
+where the distance-prior is weighted with the attention score for both interaction and selection. is inner product, is Hadamard product, and is the channel dimension.
 Coarse-to-Fine Selection.
-To capture the interaction information from coarse to fine, we stack dual-interaction layers in a cascaded manner, where a top-K operation is appended between each two consecutive layers, thus the interactive objects can be gradually selected for latter prediction and planning usages.
+To capture the interaction information from coarse to fine, we stack dual-interaction layers in a cascaded manner, where a top- operation is appended between each two consecutive layers, thus the interactive objects can be gradually selected for latter prediction and planning usages.
 We claim that only a few interactive objects need to be considered for motion prediction, which are enough yet efficient for ego-centric path planning, instead of all detected agents existing in the driving scene.
+is set to 3 and empirically.
 3.5 Iterative Motion Planner
 As shown in Fig.
-2, the iterative motion planner is designed to conduct motion prediction for both interactive agents and ego-vehicle, and then optimize the proposal ego-trajectory with both safety and kinematic constrains iteratively.
+2, the iterative motion planner is designed to conduct motion prediction for both interactive agents and ego-vehicle, and then optimize the ego-trajectory with both safety and kinematic constrains iteratively.
 Joint Motion Prediction.
@@
 Another difference is that only the interactive objects (CIPV) sparsely selected in the former module are considered, instead of all detected agents in the driving scene which maybe irrelevant to the ego-vehicle planning.
-As for the joint motion decoder, we prepare three copies of ego query to indicate different driving intentions (i.e., turn left, turn right and keep forward), which are combined with to conduct agent-level self attention and agent-map cross attention respectively.
-And then we concatenate these output attended features to predict multi-modal trajectories , and classification scores , for both agents and ego-vehicle, where is the number of driving command for planning, are the mode number, are the future timestamps.
+As for the joint motion decoder, we prepare three copies of ego query to indicate different driving intentions (i.e., turn left, turn right and keep forward), which are combined with to conduct agent-level self-attention and agent-map cross attention respectively.
+And then we concatenate these enhanced features to predict multi-modal trajectories , and classification scores , for both agents and ego-vehicle, where is the number of driving command for planning, are the mode number, are the future timestamps.
 Planning Optimization.
@@
 As shown in Fig.
-3(b), ego-agent, ego-map and ego-navigator cross attentions are conducted consecutively for planning optimization, where the offsets for each future waypoint are predicted upon the proposal trajectory respectively with several planning constraints proposed in [17] to ensure safety.
+3(b), ego-agent, ego-map and ego-navigator cross attentions are conducted consecutively for planning optimization.
+And the offsets for each future waypoint are predicted upon the proposal trajectory respectively with several planning constraints proposed in [17] to ensure safety.
 Iterative Refinement.
-To further promote the stability and performance of the whole end-to-end system, an additional iterative refinement strategy is proposed to continuously update the reference line and distance map with refined ego trajectory as illustrated in Fig.
-2, thus ensuring the interaction quality and selection accuracy of interactive queries.
+To ensure the interaction quality and selection accuracy of interactive queries, an additional iterative refinement strategy is proposed to continuously update the reference line and distance map with refined ego trajectory as illustrated in Fig.
+2.
 3.6 Uncertainty Denoising
@@
 Under this circumstance, we propose a two-level uncertainty modeling strategy to further stabilize the whole framework.
-On one hand, position-level diffusion process is performed on ground-truth boxes of interactive agents for additional trajectory prediction of noisy agents:
+On one hand, position-level diffusion process is performed on Top- boxes of interactive agents for additional trajectory prediction of noisy agents:
 which are equipped with groups of random noises following uniform distributions.
@@
 On the other hand, trajectory-level denoising process is also introduced for robust offset prediction of proposal trajectory of ego-vehicle in the planning optimization stage.
-Different from the position diffusion of detection or motion query described above, we apply the random noise to trajectory offsets of ego-vehicle , where depends on the Final Displacement (FD) of ground-truth ego future trajectory.
+Different from the position diffusion of agent query on the purpose of detection and motion, we apply the random noise to the ground-truth trajectory of ego-vehicle:
+where indicate noisy trajectory proposals and the noise scale depends on the Final Displacement (FD) of .
 | Protocol | Method | Backbone | L2 (m) | Collision (%) | Latency (ms) | FPS | ||||||
 | 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
-| ST-P3 Metrics | ST-P3 [10] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 |
-| OccNet [35] | ResNet101-DCN | 1.29 | 2.13 | 2.99 | 2.13 | 0.21 | 0.59 | 1.37 | 0.72 | - | - | |
-| FusionAD∗ [39] | ResNet101+SECOND [37] | - | - | - | 1.03 | 0.25 | 0.13 | 0.25 | 0.21 | - | - | |
+| ST-P3 Metrics | FF∗ [10] | - | 0.55 | 1.20 | 2.54 | 1.43 | 0.06 | 0.17 | 1.07 | 0.43 | - | - |
+| EO∗ [18] | - | 0.67 | 1.36 | 2.78 | 1.60 | 0.04 | 0.09 | 0.88 | 0.33 | - | - | |
+| ST-P3 [11] | EfficientNet-b4 | 1.33 | 2.11 | 2.90 | 2.11 | 0.23 | 0.62 | 1.27 | 0.71 | 628.3 | 1.6 | |
 | VAD-Base [17] | ResNet50 | 0.41 | 0.70 | 1.05 | 0.72 | 0.07 | 0.17 | 0.41 | 0.22 | 224.3 | 4.5 | |
 | SparseDrive-S [39]† | ResNet50 | 0.30 | 0.58 | 0.95 | 0.61 | 0.47 | 0.47 | 0.69 | 0.54 | 111.1 | 9.0 | |
-| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 | 67.7 | 14.8 | |
+| EgoFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 | 67.7 | 14.8 | |
 | SparseDrive Metrics‡ | UniAD [12]† | ResNet101-DCN | 0.45 | 0.70 | 1.04 | 0.73 | 0.62 | 0.58 | 0.63 | 0.61 | 555.6 | 1.8 |
@@
 | SparseDrive-B [39] | ResNet101 | 0.29 | 0.55 | 0.91 | 0.58 | 0.01 | 0.02 | 0.13 | 0.06 | 137.0 | 7.3 | |
-| DiFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
-| DiFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 93.7 | 10.7 | |
-| DiFSD-B (PV) | ResNet101 | 0.15 | 0.30 | 0.54 | 0.32 | 0.00 | 0.04 | 0.15 | 0.06 | 119.6 | 8.4 | |
+| EgoFSD-S (BEV) | ResNet50 | 0.16 | 0.33 | 0.59 | 0.35 | 0.03 | 0.07 | 0.21 | 0.10 | 67.7 | 14.8 | |
+| EgoFSD-S (PV) | ResNet50 | 0.15 | 0.31 | 0.56 | 0.33 | 0.00 | 0.06 | 0.19 | 0.08 | 59.8 | 16.7 | |
+| EgoFSD-B (PV) | ResNet101 | 0.12 | 0.28 | 0.52 | 0.30 | 0.00 | 0.04 | 0.15 | 0.06 | 80.5 | 12.4 | |
 3.7 End-to-End Learning
@@
 Meanwhile, vectorized planning constrains identified in [17] such as collision, overstepping and direction are also included in for regularization.
+Besides, the weight terms of different losses are adjusted empirically to ensure the same magnitude.
 is the number of motion planning stages.
@@
 4.2 Implementation Details
-DiFSD plans a 3 future ego-trajectory with 2 history information as input.
-Our DiFSD has two variants, namely DiFSD-S and DiFSD-B.
-As for DiFSD-S, both perspective-view perception version DiFSD-S (PV) and BEV perception version DiFSD-S (BEV) are all implemented for comparison.
+EgoFSD plans a 3 seconds future ego-trajectory with 2 seconds history information as input, which has two variants, namely EgoFSD-S and EgoFSD-B.
+As for EgoFSD-S, both Perspective-View version and BEV version are all implemented.
 ResNet50 [8] is adopted as the default backbone network for visual encoding.
 The perception range is set to 6030 longitudinally and laterally.
-Input image size of DiFSD-S is resized to .
-For DiFSD-S (BEV), the default number of BEV query, map query, agent query is , and 300, respectively.
-For DiFSD-S (PV), is 6, is 900 and is 100 respectively.
+Input image size of EgoFSD-S is resized to .
+For EgoFSD-S (BEV), is 3, and the number of BEV query, map query, agent query are , and 300, respectively.
+For EgoFSD-S (PV), is 900 and is 100 respectively.
 Each map element contains 20 map points.
@@
 is set to 3.
-DiFSD-B has larger input image resolution ( and backbone network (ResNet101).
-We use AdamW [30] optimizer and Cosine Annealing [29] scheduler to train DiFSD with weight decay 0.01 and initial learning rate -4.
-DiFSD is trained for 60 epochs and 4 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A100 GPUs with batch size 8 empirically.
+EgoFSD-B has deeper network (ResNet101, = 6) and larger image resolution (.
+We use AdamW [33] optimizer and Cosine Annealing [32] scheduler to train EgoFSD with weight decay 0.01 and initial learning rate -4.
+EgoFSD is trained for 60 epochs and 6 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A800 GPUs with total batch size 32 empirically.
+| Method | Open-loop Metric | Closed-loop Metrics | ||
+| Avg. L2 (m) | Driving Score | Success Rate (%) | Efficiency | |
+| AD-MLP [49] | 3.64 | 18.05 | 0.00 | 48.45 |
+| UniAD-Tiny [12] | 0.80 | 40.73 | 13.18 | 123.92 |
+| UniAD-Base [12] | 0.73 | 45.81 | 16.36 | 129.21 |
+| VAD [17] | 0.91 | 42.35 | 15.00 | 157.94 |
+| GenAD [52] | - | 44.81 | 15.90 | - |
+| MomAD [38] | 0.82 | 44.54 | 16.71 | 170.21 |
+| EgoFSD-S | 0.70 | 52.02 | 21.00 | 178.30 |
+| EgoFSD-B | 0.66 | 60.39 | 31.78 | 180.63 |
 4.3 Main Results
 Open-Loop Planning Evaluation.
-As show in Tab.
-1, DiFSD shows great advantages in both performance and efficiency compared with previous works.
-On one hand, DiFSD-S achieves the minimum L2 error even with lightweight visual backbone and inferior BEV perception encoder.
-Specifically, compared with BEVFormer-based end-to-end methods [11, 17], DiFSD-S (BEV) reduces the average L2 error by a great margin (0.38 and 0.37, separately), while significantly reducing the average collision rates by 84% and 52% respectively.
-Equipped with deeper visual backbone and advanced sparse detectors from Perspective View (PV), the average L2 error and collision rates can be further reduced to 0.32 and to 0.06 respectively.
-Notably, we are the first to achieve 0 collision rate on 1.
+As shown in Tab.
+1, EgoFSD demonstrates significant advantages in both performance and efficiency compared to previous works.
+On one hand, EgoFSD-S achieves the minimum L2 error despite its lightweight visual backbone and inferior BEV perception.
+Specifically, compared with BEVFormer-based end-to-end methods [12, 17], EgoFSD-S (BEV) reduces the average L2 error by a great margin (0.38 and 0.37, separately), while significantly reducing the average collision rates by 84% and 52% respectively.
+Equipped with deeper visual backbone and advanced sparse detectors from Perspective View (PV), the average L2 error and collision rates can be further reduced to 0.30 and to 0.06 respectively.
 On the other hand, benefiting from the ego-centric hierarchical interaction, only sparse interactive agents (2%) are considered for motion planning.
-Hence, DiFSD-S can achieve great efficiency with 14.8 FPS, and faster than UniAD [11] and VAD [17] respectively.
+Hence, EgoFSD-S can achieve great efficiency with 16.7 FPS, and faster than UniAD [12] and VAD [17] respectively.
 Closed-Loop Planning Evaluation.
@@
 UniAD [12] has a lower L2 error compared to VAD [17] but with worse closed-loop planning performance as discussed in [23].
-Notably, DiFSD-S achieves both the lowest L2 error and best closed-loop performance with great efficiency, showcasing the superiority and generalizability of our proposed method.
+Notably, EgoFSD-S achieves both the lowest L2 error and best closed-loop performance with great efficiency, showcasing the superiority and generalizability of our proposed method.
 4.4 Ablation Study
-We conduct extensive experiments to study the effectiveness of our DiFSD. We use DiFSD-S as default for ablation. More experiments are provided in the Appendix.
+We conduct extensive experiments to study the effectiveness of our EgoFSD. We use EgoFSD-S as default for ablation. More experiments are provided in the Appendix.
 Necessity of Geometric Prior.
@@
 While using the ego-centric cross attention, only 2% of surrounding queries are enough for achieving convincing planning performance.
-Besides, introducing the geometric prior through attention can dramatically reduce the L2 error and collision rate by 8% and 42% respectively.
+Besides, introducing the geometric prior through attention can dramatically reduce the collision rate by 42% especially, thanks to the superior interactive planning.
 Meanwhile, when utilizing the Ground-Truth geometric score for upper-limit evaluation, we can obtain the extremely excellent planning performance (0.23 average L2 error).
-Undoubtedly, the proposed ego-centric query selector equipped with geometric attention is nontrivial for motion planner.
+The proposed ego-centric query selector equipped with geometric attention is nontrivial for motion planner.
 Effect of designs in Hierarchical Interaction.
@@
 4 shows the effectiveness of our elaborate designs in the hierarchical interaction module, which contains three main designs such as Dual Interaction (DI), Geometric Attention (GA) and Coarse-to-Fine Selection (CFS).
-DI models both ego-centric and object-centric interactions respectively, which improves the planning performance greatly as expected.
+DI models both ego-centric and object-centric interactions respectively, which ensures the query selection quality for interactive planning.
 GA facilitates the query selection process as discussed in Tab.
-3, which reduces the collision rate by a great margin (42%).
-And CFS contributes to the interaction modeling quality through hierarchical receptive fields from global to local.
+3, which reduces the collision rate obviously owing to the specialized attention on CIPV / CIPS.
+And CFS contributes to the multi-granularity interaction modeling through hierarchical receptive fields from global to local.
 All of these three designs combined together can achieve overall convincing planning performance.
@@
 | Attn (5%) | ✗ | 0.16 | 0.34 | 0.63 | 0.38 | 0.07 | 0.09 | 0.31 | 0.16 |
-| Attn (2%) | ✗ | 0.16 | 0.34 | 0.61 | 0.37 | 0.08 | 0.11 | 0.27 | 0.15 |
+| Attn (2%) | ✗ | 0.16 | 0.34 | 0.61 | 0.37 | 0.06 | 0.08 | 0.24 | 0.12 |
 | Attn (2%) | Random | 0.17 | 0.36 | 0.67 | 0.40 | 0.07 | 0.10 | 0.34 | 0.17 |
 | Attn (2%) | GroundTruth | 0.14 | 0.23 | 0.33 | 0.23 | 0.07 | 0.08 | 0.10 | 0.07 |
-| Attn (2%) | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
+| Attn (2%) | ✓ | 0.16 | 0.33 | 0.59 | 0.35 (-5%) | 0.00 | 0.04 | 0.18 | 0.07 (-42%) |
 | DI | GA | CFS | Planning L2 (m) | Planning Coll. (%) | ||||||
 | 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||
-| ✗ | ✓ | ✓ | 0.18 | 0.35 | 0.62 | 0.38 | 0.09 | 0.12 | 0.23 | 0.14 |
-| ✓ | ✗ | ✓ | 0.16 | 0.34 | 0.61 | 0.37 | 0.08 | 0.11 | 0.27 | 0.15 |
-| ✓ | ✓ | ✗ | 0.16 | 0.33 | 0.59 | 0.36 | 0.09 | 0.11 | 0.25 | 0.15 |
+| ✗ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.36 | 0.01 | 0.08 | 0.23 | 0.11 |
+| ✓ | ✗ | ✓ | 0.16 | 0.34 | 0.61 | 0.37 | 0.06 | 0.08 | 0.24 | 0.12 |
+| ✓ | ✓ | ✗ | 0.19 | 0.37 | 0.64 | 0.40 | 0.09 | 0.12 | 0.23 | 0.14 |
 | ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
+| ID | JMP | PO | IR | UD | Planning L2 (m) | Planning Coll. (%) | ||||||
+| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
+| 1 | ✓ | ✗ | ✗ | ✓ | 0.23 | 0.48 | 0.83 | 0.51 | 0.08 | 0.13 | 0.35 | 0.18 |
+| 2 | ✓ | ✓ | ✗ | ✓ | 0.16 | 0.33 | 0.61 | 0.37 | 0.01 | 0.08 | 0.23 | 0.11 |
+| 3 | ✓ | ✓ | ✓ | ✗ | 0.16 | 0.34 | 0.64 | 0.38 | 0.07 | 0.07 | 0.17 | 0.10 |
+| 4 | ✓ | ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
 Effect of designs in Motion Planner.
-As for motion planner in DiFSD, Joint Motion Prediction (JMP), Planning Optimization (PO) as well as Iterative Refinement (IR) makes up the planning pipeline of ego-vehicle.
+As for motion planner in EgoFSD, Joint Motion Prediction (JMP), Planning Optimization (PO) as well as Iterative Refinement (IR) makes up the planning pipeline of ego-vehicle.
 Besides, Uncertain Denoising (UD) contributes to the system stability and training convergence.
@@
 6.
-We can observe that our DiFSD can obtain superior planning performance with one additional refinement stage (36.3% collision rate reduction), which becomes saturated when introducing more stages.
+We can observe that our EgoFSD can obtain superior planning performance with one additional refinement stage (36.3% collision rate reduction), which becomes saturated when introducing more stages.
 Hence, two-stage interacted motion planner is enough for achieving convincing results.
@@
 7, motion diffusion can improve the prediction stability with uncertain agent positions, while the planning denoising can also strengthen the trajectory regression precision of ego-vehicle.
-| ID | JMP | PO | IR | UD | Planning L2 (m) | Planning Coll. (%) | ||||||
-| 1s | 2s | 3s | Avg. | 1s | 2s | 3s | Avg. | |||||
-| 1 | ✓ | ✗ | ✗ | ✓ | 0.23 | 0.48 | 0.83 | 0.51 | 0.08 | 0.13 | 0.35 | 0.18 |
-| 2 | ✓ | ✓ | ✗ | ✓ | 0.16 | 0.33 | 0.61 | 0.37 | 0.01 | 0.08 | 0.23 | 0.11 |
-| 3 | ✓ | ✓ | ✓ | ✗ | 0.16 | 0.34 | 0.64 | 0.38 | 0.07 | 0.07 | 0.17 | 0.10 |
-| 4 | ✓ | ✓ | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
 | Number of Stages | Planning L2 (m) | Planning Coll. (%) | ||||||
@@
 | 4 | 0.16 | 0.33 | 0.61 | 0.36 | 0.00 | 0.04 | 0.20 | 0.08 |
-Runtime of each module.
-We evaluate the modular runtime of DiFSD-S with different perception paradigms as shown in Tab.
-8.
-Generally, visual backbone and perception encoder occupy the most of the runtime for feature extraction and scene understanding.
-Hierarchical interaction also takes a significant part for interaction modeling and interactive query selection.
-Thanks to the ego-centric interaction and selection module, the joint motion planner only consumes 7.9ms to plan the future ego-trajectory.
-Notably, though PV-based perception module [24] achieves superior detection performance, it consumes more computation cost due to the repeated query projection, deformable feature aggregation as well as the symmetric sparse perception architecture for unified object detection and online mapping, without a shared BEV feature as [17, 11].
-4.5 Qualitative Results
-We visualize the motion trajectories of sparse interactive agents as well as planning results of DiFSD as illustrated in Fig.
-5.
-Both surrounding camera images and prediction results on BEV are provided accordingly.
-Besides, we also project the planning trajectories to the front-view camera image.
-Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of DiFSD motivation.
-DiFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
 | Position Diffusion | Trajectory Denoising | Planning L2 (m) | Planning Coll. (%) | ||||||
@@
 | ✓ | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
-| Module | Bird’s-Eye View (BEV) Perception | Perspective-View (PV) Perception | ||
-| Latency (ms) | Proportion (%) | Latency (ms) | Proportion (%) | |
-| Backbone | 8.4 | 12.4 | 11.0 | 11.8 |
-| Perception Encoder | 34.4 | 50.8 | 54.9 | 58.6 |
-| Hierarchical Interaction | 17.0 | 25.1 | 19.9 | 21.2 |
-| Motion Prediction | 4.5 | 6.6 | 4.5 | 4.8 |
-| Planning Optimization | 3.4 | 5.1 | 3.4 | 3.6 |
-| Total | 67.7 | 100 | 93.7 | 100 |
+Runtime of each module.
+We evaluate the modular runtime of EgoFSD-S with Perspective-View perception paradigms as shown in Tab.
+8.
+Generally, visual backbone and perception decoder occupy the most of the runtime (57.7%) for feature extraction and scene understanding.
+Hierarchical interaction also takes a significant part for interactive query selection.
+Thanks to the ego-centric interaction and selection module, the joint motion planner only consumes 7.5ms to plan the future ego-trajectory.
+4.5 Qualitative Results
+We visualize the motion trajectories of sparse interactive agents as well as planning results of EgoFSD as illustrated in Fig.
+A3.
+Both surrounding camera images and prediction results on BEV are provided accordingly.
+Besides, we also project the planning trajectories to the front-view camera image.
+Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of EgoFSD motivation.
+EgoFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
+| Module | #Param (M) | Latency (ms) | Proportion (%) |
+| Backbone | 26.7 | 4.7 | 7.9 |
+| Sparse Perception | 11.2 | 29.8 | 49.8 |
+| Hierarchical Interaction | 21.3 | 17.8 | 29.8 |
+| Joint Motion Prediction | 4.9 | 4.1 | 6.9 |
+| Planning Optimization | 1.8 | 3.4 | 5.6 |
+| Total | 65.9 | 59.8 | 100.0 |
 5 Conclusion
-In this paper, we propose a fully sparse paradigm for end-to-end self-driving in an ego-centric manner, termed as DiFSD.
-DiFSD revisits the human driving behavior and conducts hierarchical interaction based on sparse representation and perception results.
-Only interactive agents are considered for joint motion prediction with the ego-vehicle.
-Iterative planning optimization strategy contributes to the driving safety with high efficiency.
+In this paper, we propose a Fully Sparse Paradigm for end-to-end self-driving in an Ego-Centric manner, termed as EgoFSD.
+EgoFSD conducts hierarchical interaction based on sparse representation and perception results.
+Only interactive agents are considered for joint motion prediction including the ego-vehicle.
+Iterative planning optimization strategy contributes to the driving safety with interactive decision.
 Besides, uncertainty modeling is conducted to improve the stability of end-to-end system.
-Extensive ablations and comparisons reveal the superiority and great potential of our ego-centric fully sparse paradigm for future research.
+Extensive ablations and comparisons reveal the superiority and great potential of our ego-centric fully sparse paradigm.
 Supplementary Material
@@
 Perception.
-The evaluation for detection and tracking follows standard evaluation protocols [1].
+The evaluation for detection and online mapping follows standard evaluation protocols [1].
 For detection, we use mean Average Precision (mAP), mean Average Error of Translation (mATE), Scale (mASE), Orientation (mAOE), Velocity (mAVE), Attribute (mAAE) and nuScenes Detection Score (NDS) to evaluate the model performance.
 For online mapping, we calculate the Average Precision (AP) of three map classes: lane divider, pedestrian crossing and road boundary, then average across all classes to get mean Average Precision (mAP).
-Planning.
+Open-Loop Planning.
 We adopt commonly used L2 error and collision rate to evaluate the planning performance.
@@
 We reproduce the planning results on our benchmark with official checkpoints [12, 17] for a fair comparison.
-Appendix B More Ablation Study
+Closed-Loop Planning.
+We use the official 220 routes and official metrics of Bench2Drive [15] for evaluation.
+The Driving Score (DS) is defined as the product of Route Completion and Infraction Score, capturing both task completion and rule adherence.
+The Success Rate (SR) measures the percentage of routes completed successfully within the allocated time.
+A route is deemed successful if the ego vehicle reaches its destination without any traffic rule violation.
+Efficiency quantifies the vehicle’s velocity relative to surrounding traffic, encouraging progressiveness without aggression.
+Appendix B More Details
+Problem Formulation.
+Given multi-view camera image sequence can be denoted as , where is the number of camera views and indicates the temporal length till current timestep respectively.
+Annotation of input for end-to-end planning is composed by a set of future waypoints of the ego-vehicle , where = 3 seconds (2Hz) is the planning time horizon, and (, ) is the BEV location transformed to the ego-vehicle coordinate system at current timestep .
+Meanwhile, driving command as well as ego-status is also provided.
+Annotation set is used during training.
+During prediction, the planned trajectory of ego-vehicle should fit the annotation with minimum L2 errors and collision rate with surrounding agents.
+Data Augmentation.
+During training phase, we conduct a global data augmentation strategy to improve the model stability.
+Specifically, random rotation, translation and flipping (Y-axis) are applied to both bounding boxes of agent / map instances and future trajectories of agents / ego-vehicle respectively.
+Meanwhile, we adjust the camera extrinsic parameters accordingly to ensure spatial alignment between the surround-view images and the noise-perturbed targets.
+Appendix C More Ablation Study
 Necessity and Order of Object Selection.
@@
 And both of agent and map interactions are conducted in a cascaded order is inferior than the parallel manner, where the updated ego query from parallel outputs are concatenated for joint motion prediction.
+Effect of Interactive Score Fusion.
+During the ego-centric query selection, both geometric and classification scores are considered to ensure that the selected closest in-path queries are true positive agents or maps, which are adopted for motion planner.
+Tab.
+2 shows the effect of three types of scores used for query ranking, namely attention, geometry and confidence scores.
+As Eq.
+3, interactive score obtained by multiplying these three scores can achieve the best selection quality and planning performance.
+without confidence score fails to distinguish between background and foreground queries, resulting in inferior performance.
 | Agent Selection | Map Selection | Cascade | Parallel | Planning L2 (m) | Planning Coll. (%) | ||||||
@@
 | ✓ | ✓ | - | ✓ | 0.16 | 0.33 | 0.59 | 0.35 | 0.00 | 0.04 | 0.18 | 0.07 |
-Effect of Interactive Score Fusion.
-During the ego-centric query selection, both geometric and classification scores are considered to ensure that the selected closest in-path queries are true positive agents or maps, which are adopted for motion planner.
-Tab.
-2 shows the effect of three types of scores used for query ranking, namely attention, geometry and confidence scores.
-As Eq.
-2, interactive score obtained by multiplying these three scores can achieve the best selection quality and planning performance.
-without confidence score fails to distinguish between background and foreground queries, resulting in inferior performance.
 | Attention Score | Geometric Score | Classification Score | Planning L2 (m) | Planning Coll. (%) | ||||||
@@
 | UniAD [12] | ResNet101-DCN | ✓ | 38.0 | 49.8 |
-| VAD† [17] | ResNet50 | ✓ | 27.3 | 39.7 |
+| VAD-Base† [17] | ResNet50 | ✓ | 31.3 | 43.6 |
 | SparseDrive-S [39] | ResNet50 | ✗ | 41.8 | 52.5 |
 | SparseDrive-B [39] | ResNet101* | ✗ | 49.6 | 58.8 |
-| DiFSD-S | ResNet50 | ✓ | 32.8 | 45.8 |
-| DiFSD-S | ResNet50 | ✗ | 41.0 | 52.8 |
-| DiFSD-B | ResNet101* | ✗ | 49.6 | 58.9 |
+| EgoFSD-S (BEV) | ResNet50 | ✓ | 32.8 | 45.8 |
+| EgoFSD-S (PV) | ResNet50 | ✗ | 41.0 | 52.8 |
+| EgoFSD-B (PV) | ResNet101* | ✗ | 49.6 | 58.9 |
 | Method | APped | APdivider | APboundary | mAP |
@@
 | MapTR [25] | 56.2 | 59.8 | 60.1 | 58.7 |
-| VAD† [17] | 40.6 | 51.5 | 50.6 | 47.6 |
+| VAD-Base† [17] | 42.5 | 50.5 | 49.8 | 47.6 |
 | SparseDrive-S [39] | 49.9 | 57.0 | 58.4 | 55.1 |
 | SparseDrive-B [39] | 53.2 | 56.3 | 59.1 | 56.2 |
-| DiFSD-S (BEV) | 46.7 | 54.3 | 56.0 | 52.3 |
-| DiFSD-S (PV) | 54.9 | 55.7 | 57.3 | 56.0 |
-| DiFSD-B (PV) | 52.3 | 58.2 | 59.3 | 56.6 |
-Appendix C Analysis & Discussion
-The GroundTruth future state distribution of ego-vehicle on nuScenes validation set is illustrated in Fig.
-A1, which is calculated with fixed time interval (1) between consecutive predicted waypoints.
-And we also compare the output ego-state distribution of different popular end-to-end methods based on planned trajectories respectively, as show in Fig.
+| EgoFSD-S (BEV) | 46.7 | 54.3 | 56.0 | 52.3 |
+| EgoFSD-S (PV) | 54.9 | 55.7 | 57.3 | 56.0 |
+| EgoFSD-B (PV) | 52.3 | 58.2 | 59.3 | 56.6 |
+Appendix D Effectiveness Analysis
+The Ground-Truth future state distribution of ego-vehicle on nuScenes validation set is illustrated in Fig.
+A1, which is calculated with fixed time interval (1 second) between consecutive output future waypoints.
+And we also compare the output ego-state distribution of different popular E2E methods based on planned trajectories respectively, as show in Fig.
 A2.
-We can observe that without ego-centric design, the optimized end-to-end model is unable to handle various emergencies appearing in the driving scenarios, where the absolute values of and are larger than normal situations.
-Under this circumstance, the output planned trajectories cannot conform to the expert routes as expected.
-However, our DiFSD performs consistently better in planning the future ego states with variable speed and acceleration, owing to the ego-centric hierarchical interaction and selection mechanism, thus the iterative motion planner can focus on the interactive agents rather than irrelevant objects.
-Appendix D Visualization
+We can observe that without ego-centric design, existing end-to-end models are incapable of handling various emergencies appearing in the driving scenarios, where the absolute values of and are larger than normal situations (i.e., emergent braking, acceleration, turning or obstacle avoiance).
+Under this circumstance, the output planned trajectories cannot conform to the recorded expert demonstrations as expected.
+However, our EgoFSD performs consistently better in planning the future trajectory with dynamic and adaptive ego states of variable speed and acceleration.
+Owing to the ego-centric hierarchical interaction and selection mechanism, the iterative motion planner can focus on the interactive agents rather than general objects.
+Appendix E More Discussion
+Practicality and Generalizability of Ego-Centric Design.
+We empirically validate the necessity and effectiveness of our proposed ego-centric query selection and interactive planning with the help of geometric prior in Sec.
+4.4.
+Moreover, we claim that the usage of geometric prior through interactive score fusion is general and practical for data-driven E2E-AD training, which considers both semantic relevance of driving context and human-like driving attention distribution according to current driving situation and future goal waypoint provided by the route planner.
+Extensive experiments conducted on both popular nuScenes [1] and large-scale challenging Bench2Drive [15] benchmark demonstrate the generalizability and significant improvement of computational efficiency.
+Comparison to Diffusion-based Planners.
+We notice that recent works [26, 41] also explore diffusion model for end-to-end autonomous driving application.
+DiffusionDrive [26] learns denoising from anchored Gaussian distribution to the multi-mode driving action distribution with a truncated diffusion policy.
+DiffAD [41] treats E2E-AD as a conditional image generation task, which adopts diffusion model to predict noise from the noisy latent BEV image.
+Differently, we conduct a two-level data augmentation strategy to enrich the training samples including position diffusion of detected agents and trajectory diffusion of proposal ego trajectory.
+In this way, the joint motion prediction result can be more robust even with uncertain agent positions.
+And the offset prediction of ego trajectory can be more accurate in the planning optimization stage.
+Extensive ablations demonstrate the effectiveness of our uncertainty denoising design in improving the performance and training stability of EgoFSD.
+Comparison to other Sparse Frameworks.
+Prior to our work, some exiting methods [39, 50] are also built upon sparse representation for end-to-end autonomous driving.
+SparseDrive [39] introduces the symmetric sparse perception architecture for object detection and online mapping respectively.
+However, it consumes more computation cost due to the repeated query projection, deformable feature aggregation, without a shared BEV feature as [17, 12].
+Our EgoFSD constructs a shared 3D position-aware feature as [7] to facilitate parallel sparse query decoding for perception and planning.
+Besides, an additional core contribution of our EgoFSD lies on an ego-centric design, which aims to select the most important agent/map instances (i.e., CIPV/CIPS) for interactive modeling.
+Benefiting from the geometric prior and iterative refinement, the query selection quality can be gradually enhanced for efficient planning.
+Extensive ablations showcase the superiority of our proposed Ego-Centric Fully Sparse Paradigm.
+Appendix F Limitations and Social Impact
+Limitations.
+There still remains some limitations in our work.
+First, EgoFSD performs ego-centric query selection based on the geometric score, which is obtained through referring to the distance map using static query positions.
+And the motion information of surrounding agents are introduced into the ego-object dual interaction layer implicitly through iteratively updating the agent queries.
+How to generate better geometric score considering agent dynamics explicitly is worthy of further discussion.
+Besides, how to incorporate additional traffic signals and vision-language models into current driving system for autonomous decision making also deserves further exploration.
+Social Impact.
+EgoFSD could be easily deployed on mass-produced car chips with different limits of computing resources, and thus can be served as a plug-and-play software to assist human drivers in decision-making and safe driving.
+Appendix G Qualitative Results
+We visualize the motion trajectories of sparse interactive agents as well as planning results of EgoFSD as illustrated in Fig.
+A3.
+Both surrounding camera images and prediction results on BEV are provided accordingly.
+Besides, we also project the planning trajectories to the front-view camera image.
+Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of EgoFSD motivation.
+EgoFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
 As show in Fig.
-A3-A5, we provide more visualization results to illustrate the effectiveness of DiFSD on various driving scenarios (i.e., interactive / non-interactive scenes, lane-change / lane-keep / following scenes, overtaking / avoidance scenes, intersection scenes) under different commands (i.e., “Go Straight”, “Turn Left”, “Turn Right”).
+A4-A6, we provide more visualization results to illustrate the effectiveness of EgoFSD on various driving scenarios (i.e., interactive / non-interactive scenes, lane-change / lane-keep / following scenes, overtaking / avoidance scenes, intersection scenes) under different commands (i.e., “Go Straight”, “Turn Left”, “Turn Right”).
 Moreover, we also observe some failure cases on nuScenes validation set as illustrated in Fig.
-A6 and Fig.
-A7.
-The corresponding explanations as well as analysis are described in the captions below respectively, which mainly attribute to the irrationality of the expert (ground truth) future trajectories of ego vehicle, demonstrating the strong generalizability of our proposed DiFSD in planning the efficient and reasonable results based on the selected sparse representations of surrounding driving scenarios.
+A7 and Fig.
+A8.
+The corresponding explanations as well as analysis are described in the captions below respectively, which mainly attribute to the irrationality of the expert (ground truth) future trajectories of ego vehicle, demonstrating the strong generalizability of our proposed EgoFSD in planning the efficient and reasonable results based on the selected sparse representations of surrounding driving scenarios.